# Egocentric Scene-aware Human Trajectory Prediction

### v1

@@ Abstract
Wearable collaborative robots stand to assist human wearers who need fall prevention assistance or wear exoskeletons. Such a robot needs to be able to predict the ego motion of the wearer based on egocentric vision and the surrounding scene. In this work, we leveraged body-mounted cameras and sensors to anticipate the trajectory of human wearers through complex surroundings. To facilitate research in ego-motion prediction, we have collected a comprehensive walking scene navigation dataset centered on the user’s perspective. We present a method to predict human motion conditioning on the surrounding static scene. Our method leverages a diffusion model to produce a distribution of potential future trajectories, taking into account the user’s observation of the environment. We introduce a compact representation to encode the user’s visual memory of the surroundings, as well as an efficient sample-generating technique to speed up real-time inference of a diffusion model. We ablate our model and compare it to baselines, and results show that our model outperforms existing methods on key metrics of collision avoidance and trajectory mode coverage.
Keywords:
Egocentric Trajectory Prediction Diffusion Model Scene Understanding
Figure 1: Different red ribbon illustrates different possible modes in the scene and the size of the ribbon denotes the likelihood.
1 Introduction
@@
The integration of autonomous systems into human-centric environments, particularly through collaborative robotics, necessitates acute awareness and prediction of human motion. These systems, whether external like autonomous vehicles and mobile robots or wearable like exoskeletons, require precise motion prediction to ensure safety and enhance collaboration. External systems predict human trajectories from a distance to avoid collisions or facilitate physical collaboration, whereas wearable systems, mounted directly on users, rely on an egocentric perspective to predict the wearer’s motion for functional response.
@@
Predicting motion from an egocentric perspective involves deciphering the complex relationship between past movements, task semantics, and potential future actions. This task is complicated by the need to account for multi-modal outcomes, such as which path a person might choose at a divergence. Factors influencing these decisions range from high-level objectives to immediate environmental constraints like obstacles and walkable paths. A generative machine learning model can learn to predict the motion of a human by observing human data of individuals walking in different scenes. Such models are improved and more efficient when curated features are provided. Scene semantics can help provide curated features that improve generative model prediction.
@@
In this paper, we propose a generative modeling approach to predict the distribution of future trajectories of a person given the environment semantics observed from the egocentric view, as illustrated in Fig. 1. To address the under-constrained and multi-modal nature of the navigation problem, we employ a diffusion model to predict the future trajectories conditioned on the egocentric observation of the environment and the user’s past walking trajectory. We introduce a compact representation of the environment to capture a short history of visual observations from the egocentric perspective. This "visual memory" representation encodes both appearance and semantics information of the world. To achieve real-time inference, we introduce a hybrid generation technique that balances between quality and speed of sample generation from the trained diffusion model. Using our hybrid generation method, we can rapidly sample multiple future trajectories in real time, providing downstream applications a distribution of the possible paths a person might take, rather than predicting a single path.
@@
We evaluate our predictive model using real-world navigation data. We collect a comprehensive walking scene navigation dataset centered on user’s perspective in diverse indoor and outdoor scenarios. We show that our model can accurately predict a distribution of future trajectories representing human preference taking scene context into account, outperforming the baseline methods on various evaluation metrics. We also conduct ablation studies to validate the design choices in visual memory representation and hybrid generation technique. The dataset and the trained model will be made publicly available.
2 Related Work
@@
Path prediction of humans is an area of interest across robotic and computer vision applications. In this section, we highlight key works in the domains of autonomous vehicles predicting the motion of pedestrians, the prediction of humans during sports activities, and the prediction of humans in social settings leveraging social context. We also highlight recent advances in diffusion models as a method of predicting trajectories.
For autonomous vehicles (AV), predicting the motion of other vehicles and pedestrians is crucial for the AV to plan safe, informed trajectories [12]. Recent approaches focus on extracting motion prediction from demonstrations [6]. Traditionally algorithms for trajectory prediction relied on sequential prediction methods such as recurrent neural networks (RNNs), LSTMs, or GRUs [22]. However, these models show limitations when confronted with complex, multi-modal sensor data due to the hidden state size and forgetting information in long sequences. Transformer-based models provide a powerful alternative offering benefits in multi-modal sensor streams and the ability to handle long-sequence inter-dependence’s [9, 23]. For AVs, it is often convenient to reconstruct the environment from a birds-eye-view (BEV) or 2D occupancy grid provided sensory inputs [20, 16, 3]. While such representations are easy to interpret when predictions are limited to a small set of modes or choices, they introduce a “long tail" problem which is particularly impactful when the environment has less structure, presenting a need for research in representations that balance structure and flexibility in trajectory prediction [13]
Trajectory prediction of players during sports is another application with rich existing datasets where there is inherent context related to the strategy of the game being played [2]. These sporting scenarios usually take place where there are few static objects/obstacles, and playing grounds are usually flat, leaving game strategy as the sole motion intent informer [21, 2].
Trajectory prediction of humans in social settings provides motion scenarios where behaviors are conditioned on social norms and settings, and leading models (SocialGAN, SociaLSTM) take these factors into account during prediction [4, 1]. Such models have been applied to pedestrian dynamics, showcasing the potential for robots to coexist seamlessly with humans in crowded public spaces.[14] Prominent datasets facilitating these studies include UCY/ETH [7] and the Stanford Drone Dataset[14], which predominantly offer a BEV of pedestrian movement, simplifying the trajectory prediction challenge to a two-dimensional problem. This top-down perspective allows for the tracking of pedestrians through bounding boxes, offering an effective means to study and model pedestrian trajectories in a simplified setting.
The diffusion model, inspired by the physical diffusion process, was originally proposed in 2015[17] but gained significant traction with the introduction of Denoising Diffusion Probabilistic Models (DDPM) in 2020[5]. Unlike initial attempts that struggled to predict denoised inputs directly, DDPM introduced an approach that predicts the noise to be removed, marking a breakthrough in model performance. Subsequent developments, such as Latent Diffusion Models[15] for complex generation tasks, DDIM[18] for accelerated processing, and various guidance and conditioning techniques [8], have expanded the application of diffusion models to image, video, and action generation. These models excel in generating detailed and coherent outputs across a range of generative tasks, proven by many successes in image, video, and even action generation tasks.
@@
3 Method
3.1 Problem Formulation
@@ 3.2 Method Overview
Our method takes as input the past trajectory of the person recorded by an Intel Realsense T265 camera and a short history of RGBD images recorded by an outward-facing stereo camera worn by the person and predicts the future trajectories the person might take (Figure 2. The color images are semantically labeled by DINOv2 [10], while the depth images go through a preprocessing pipeline that filters out erroneously filled edges. We transform the past trajectory from a global coordinate frame to an egocentric frame defined by the +Z as opposite to the gravity vector and forward-facing direction as +X. Taking the semantically labeled images, RGB images, and processed depth images, we project and align the images to create a single panorama in the egocentric coordinate frame, referred to as "visual memory". Conditioning on the visual memory and the past ego trajectory, a diffusion model is trained to predict the future trajectory along with encoded visual observations, as auxiliary outputs. The diffusion model starts from a random noise and performs a number of denoising steps to generate the clean prediction sequence. The clean sequence is a concatenation of encoded visual memory and future ego pose trajectory. Finally, we use the VAE decoder to recover the expected future panorama.
Figure 3: Channels in Visual Memory: The visual memory integrates frames from various time steps into a single panorama. It consists of depth channel, color channel, and intensity-encoded 8-class semantic channel. 4 channels are shown in the figure.
3.3 Construction of Visual Memory
Visual memory is defined as the network retaining an ego-perspective, panoramic view representation of the surroundings. The visual memory is constructed from aligned color, depth, and semantic images. Given the camera’s intrinsic and extrinsic parameters, images in different frames can be projected to a single point cloud in the global frame. A distance-based filter is applied to remove points too far away from the current pose. The points are then projected back to the current ego frame to form a coherent representation of all the scene information gathered.
@@ 3.3 Construction of Visual Memory
Figure 4: Comparing depth frame with visual memory: A raw depth frame from stereo camera have only  90 degrees of narrow FOV and often misses important scene information. In the figure, depth frame only sees the open space in front, did not capture the stairs, the right turn path, or the wall directly to the left. The black regions are the uncovered area when stitching frames from different time step.
@@
It is important to note that the visual memory representation holds a lot more relevant information than just a single image. As shown in Fig. 4, A single image from a stereo camera only has a narrow FOV pointing directly in front. It fails to capture the objects and paths in the scene that are highly relevant to the prediction, requiring many individual frames to be sent to the prediction module, relying on the model’s capability to extract useful information. Meanwhile, the visual memory stitches the past frames together and integrates multi-modal inputs all into one single image, greatly improving the model and storage efficiency.
3.4 Hybrid Generation of Future Trajectories with Diffusion Model
@@
The prediction part of the method heavily relies on a UNet diffusion model with multi-head self-attention layers between blocks. The self-attention layers help to relate the visual memory encoding to different parts of the trajectory, thus facilitating a deeper understanding of how humans interact with environmental features. The prediction is conditioned on a single 64-dimensional encoded visual memory, and full 100 steps of 6d pose trajectory in 20 Hz. We utilize the full 100 steps as we believe in some cases the ego-centric trajectory prediction task can violate pure Markovian assumption. For example, a different past trajectory could likely affect the visibility of the scene in visual memory, causing different exploration preferences in trajectory selection. We compared it to a pure Markovian assumption in the ablation.
@@
The output of the diffusion model is a denoised sequence that contains both normalized 100 steps of ego-centric trajectory prediction and 100 steps of visual memory encoding. The visual memory encoding can be decoded to show how visual memory is expected to march forward by the model. Including the visual memory encoding in the diffusion process also provides a foundation for the model to base the predicted trajectory. In practice, we notice that the trajectory prediction samples are more stable with the visual memory prediction. In inference time, the model will take in the same condition and generate a batch of 15 different samples, forming a discrete distribution of the expected future ego trajectory modes.
@@
4 Egocentric Navigation Dataset

### v3

# EgoNav: Egocentric Scene-aware Human Trajectory Prediction

@@ Abstract
Wearable collaborative robots stand to assist human wearers who need fall prevention assistance or wear exoskeletons. Such a robot needs to be able to constantly adapt to the surrounding scene based on egocentric vision, and predict the ego motion of the wearer. In this work, we leveraged body-mounted cameras and sensors to anticipate the trajectory of human wearers through complex surroundings. To facilitate research in ego-motion prediction, we have collected a comprehensive walking scene navigation dataset centered on the user’s perspective. We then present a method to predict human motion conditioning on the surrounding static scene. Our method leverages a diffusion model to produce a distribution of potential future trajectories, taking into account the user’s observation of the environment. To that end, we introduce a compact representation to encode the user’s visual memory of the surroundings, as well as an efficient sample-generating technique to speed up real-time inference of a diffusion model. We ablate our model and compare it to baselines, and results show that our model outperforms existing methods on key metrics of collision avoidance and trajectory mode coverage.
Figure 1: Visual Representation of the Problem: Given the past in blue, and learning from single-modal ground truth in green, what are the likely future paths and their likelihood? Different red ribbon illustrates different possible modes in the scene and the size of the ribbon denotes the likelihood.
I Introduction
@@
The integration of autonomous systems into human-centric environments, particularly through collaborative robotics, necessitates acute awareness and prediction of human motion. These systems, whether as large as autonomous vehicles or wearable like exoskeletons, require precise motion prediction to ensure safety and enhance collaboration. Their successful operation relies on a foundation of continuously inferring the intention of human cooperators, through the same egocentric perspective.
Many studies [1, 2, 3, 4, 5, 6] suggest that there is a close relationship between trajectory predictor and motion planner. Traditionally, a downstream motion planning module utilizes trajectory predictions as input to generate plans within given constraints. Recently, imitation learning approaches [7] have emerged as an alternative for motion planning. These approaches learn a probability distribution of possible actions given observations, which fundamentally relies on predicting future trajectories from the agent’s perspective. This egocentric prediction is crucial for imitation learning, as it allows the agent to anticipate and replicate human-like behavior in various scenarios. Notably, both paradigms - separate planners and imitation learning methods - can significantly benefit from a robust egocentric trajectory prediction module. This paper proposes a novel egocentric human trajectory prediction model that not only advances the state of the art in prediction but also serves as a valuable component for both traditional planning systems and imitation-based approaches. In the context of embodied agents, our model represents a crucial step towards creating more natural and adaptive motion planning and control systems, regardless of the specific planning paradigm employed.
@@
Predicting motion involves deciphering the complex relationship between past movements, scene semantics, and potential future actions. This task is complicated by the need to account for multi-modal outcomes, for instance, a person could have many equally viable choices at a divergence. Factors influencing these decisions range from high-level objectives to immediate environmental constraints like obstacles and walkable paths. Previously, work like Social Force [8] attempted to solve this task through a series of non-linearly coupled equations and simplified attracting repelling dynamics. Recently learning-based solutions such as [9, 10] have leveraged the power of data to implicitly model the more complicated preferences. To further improve upon the limitations, especially in scene understanding, we propose to use semantic segmentation to add dense channel labels to the scene. So that the model can relate different behaviors and preferences back to those labels.
Many previous work [8, 11, 12] assume the environment to be fully observed, notably in a bird’s eye view (BEV). However, a fully known environment is rarely obtainable in the real world, limiting the applicability of such algorithms. The egocentric perspective has much higher availability but also brings a unique set of challenges. The first is occlusion, especially when the scene is highly three-dimensional and cluttered. Second, a stereo camera creates a huge amount of raw data, unlike bird’s eye view, it also has a limited field of view. These challenges require careful consideration that existing literature has not done a good job of tackling. That said, semantic segmentation and generative models are critical to solving this inherently multi-modal, and highly environment conditioning task.
Unlike autonomous vehicles, the walking environment is highly unstructured and has very few traffic rules to narrow down the solution space. Most work [13, 14, 15] fails to consider the environmental features and only predicts based on the past trajectory. The pinnacle of the challenge is to solve egocentric, unstructured scene navigation with interactions with other agents in the scene. However, for now, there still remains a lot of work to do without interaction.
@@
In this paper, we propose a generative modeling approach along with the environment semantics to predict the distribution of future trajectories of a person, given the egocentric view, as illustrated in Fig 1. To address the under-constrained and multi-modal nature of the navigation problem, we employ a diffusion model to predict the future trajectories conditioned on the egocentric observation of the environment and the user’s past walking trajectory. To account for numerous challenges brought by the egocentric view, we introduce a compact representation of the environment to capture a short history of visual observations from the egocentric perspective. This ”visual memory” representation encodes both the appearance and semantics information of the world. To achieve real-time inference, we introduce a hybrid generation technique that balances between quality and speed of sample generation from the trained diffusion model. Using our hybrid generation method, we can rapidly sample multiple future trajectories in real-time, providing downstream applications a distribution of the possible paths a person might take, rather than predicting a single path.
@@
We evaluate our predictive model using real-world navigation data. We collect a comprehensive walking scene navigation dataset centered on the user’s perspective in diverse indoor and outdoor scenarios. We show that our model can accurately predict a distribution of future trajectories representing human preference taking scene context into account, outperforming the baseline methods on various evaluation metrics. We also conduct ablation studies to validate the design choices in visual memory representation and hybrid generation technique. The dataset and the trained model will be made publicly available.
II Related Work
@@
Path prediction of humans is an area of interest across the robotic and computer vision fields. In this section, we highlight key works in various domains such as autonomous vehicles, pedestrians walking directions, motion during sports activities, and humans in social settings leveraging social context. We also highlight recent advances in diffusion models as a method of generating various outputs.
@@
For autonomous vehicles (AV), predicting the motion of both ego-vehicle and other traffic agents is crucial for the AV to plan safe, informed trajectories [16]. Albeit the problem setup can vary from our task, sometimes more emphasizing third-person view trajectory prediction, we can still learn from their attempts. Recent approaches focus on extracting motion prediction from demonstrations [6], and integrate directly with planner in a differentiable manner. Traditionally algorithms for trajectory prediction relied on recurrent models such as RNNs, LSTMs, or GRUs [17]. However, these models show limitations when confronted with complex, multi-modal sensor data, mainly due to the hidden state size and forgetting information in long sequences. Transformer-based models provide a powerful alternative offering benefits in multi-modal sensor streams and the ability to handle long-sequence inter-dependence’s [18, 19]. For AVs, with numerous guiding traffic rules, it is often convenient to reconstruct the environment from a birds-eye-view (BEV) or 2D occupancy grid to fuse sensory inputs [20, 21, 22]. While such representations are easy to interpret and predictions are limited to a small set of modes or choices, they introduce a ”long tail” problem which is particularly troublesome when the structured environment assumption crumbles, presenting a need for research in representations that balance structure and flexibility in trajectory prediction.[23]
Trajectory prediction of players during sports is another field of interest [24]. The task in this field is more focused on game strategy and interactions, and less on the environmental features. Thus sporting scenarios usually take place where there are a few, if any, static objects/obstacles, and playing grounds are usually flat, leaving game strategy as the sole motion intent informer [25, 24].
Trajectory prediction of humans in walking settings is the domain this work most interested in. As mentioned in Sec.I, initial attempts usually ignore the environmental features (or assume them to be fully known), and operate on a highly abstracted level. For example Social Force [8] models all agents as a particle and moves based on various attractive and repelling forces. More recent examples are Social GAN [26], and Social LSTM [27], which assume an open space with only pedestrians. This series of work including some more recent ones [13, 14, 15] have taken the interaction modeling to its extreme, they operate by analyzing only past trajectories and predicts future paths where behaviors are purely conditioned on social norms and settings. So far, such models have achieved some promising results in terms of prediction quality running offline. However, the complete overlook of environmental interactions and overly ideal assumptions have hindered the real-world applicability of such a model. Prominent datasets facilitating these studies include UCY/ETH [28] and the Stanford Drone Dataset[29], which predominantly offer a BEV of pedestrian movement, simplifying the trajectory prediction challenge to a two-dimensional problem. This top-down perspective allows for the tracking of pedestrians through bounding boxes, offering an effective means to study and model pedestrian trajectories in a simplified setting.
Realizing that those problems have been preventing the deployment in the real world, the latest work has focused more on the perception side. Works in autonomous vehicle community [11, 12, 1, 2] have first adopted synthesized BEV as a structured fusion of environmental cues. However, the less structured walking scenarios demand a more flexible representation than BEV. So far, there are some early work [10, 30] using first-person views RGB(D) as input. More recent work has attempted more sophisticated fusing methods such as depth panorama [9], and pose estimation [31] for more informative representation. We believe there is still potential yet to be leveraged in this direction.
The diffusion model, inspired by the physical diffusion process, was originally proposed in 2015 [32] but gained significant traction with the introduction of Denoising Diffusion Probabilistic Models (DDPM) in 2020 [33]. Unlike initial attempts that struggled to predict denoised inputs directly, DDPM introduced an approach that predicts the noise to be removed, marking a breakthrough in model performance. Subsequent developments, such as Latent Diffusion Models [34] for complex generation tasks, DDIM [35] for accelerated processing, and various guidance and conditioning techniques [36], have expanded the application of diffusion models to image, video, and action generation. These models excel in generating detailed and coherent outputs across a range of generative tasks, proven by many successes in image, video, and even action-generation tasks.
III Method
Figure 3: Channels in Visual Memory: The visual memory integrates frames from various time steps into a single panorama. It consists of a depth channel, color channel, and intensity-encoded 8-class semantic channel. 4 channels are shown in the figure.
III-A Problem Formulation
@@ III-B Method Overview
Our method takes as input the past trajectory of the person recorded by an Intel Realsense T265 camera and a short history of RGBD images recorded by an outward-facing stereo camera worn by the person and predicts the future trajectories the person might take (Figure 2. The color images are semantically labeled by DINOv2 [37], while the depth images go through a preprocessing pipeline that filters out erroneously filled edges. We transform the past trajectory from a global coordinate frame to an egocentric frame defined by the +Z as opposite to the gravity vector and forward-facing direction as +X. Taking the semantically labeled images, RGB images, and processed depth images, we project and align the images to create a single panorama in the egocentric coordinate frame, referred to as ”visual memory”. Conditioning on the visual memory and the past ego trajectory, a diffusion model is trained to predict the future trajectory along with encoded visual observations, as auxiliary outputs. The diffusion model starts from a random noise and performs a number of denoising steps to generate the clean prediction sequence. The clean sequence is a concatenation of encoded visual memory and future ego pose trajectory. Finally, we use the VAE decoder to recover the expected future panorama.
@@
Figure 4: Comparing depth frame with visual memory: A raw depth frame from a stereo camera has only  90 degrees of narrow FOV and often misses important scene information. In the figure, the depth frame only sees the open space in front and does not capture the stairs, the right turn path, or the wall directly to the left. The black regions are the undiscovered areas when stitching frames from different time steps.
III-C Construction of Visual Memory
Visual memory is defined as an ego-perspective, panoramic view representation of the surroundings. The visual memory is constructed from aligned color, depth, and semantic images (Shown in Fig 3). Given the camera’s intrinsic and extrinsic parameters, images in different frames can be projected to a single point cloud in the global frame. A distance-based filter is applied to remove points too far away from the current pose. The points are then projected back to the current ego frame to form a coherent representation of all the scene information gathered.
@@ III-C Construction of Visual Memory
It is important to note that the visual memory representation holds a lot more relevant information than just a single image. As shown in Fig 4, A single image from a stereo camera only has a narrow FOV pointing directly in front. It fails to capture the objects and paths in the scene that are highly relevant to the prediction, requiring many individual frames to be sent to the prediction module, relying on the model’s capability to extract useful information. Meanwhile, the visual memory stitches the past frames together and integrates multi-modal inputs all into one single image, greatly improving the model and storage efficiency.
III-D Diffusion Pipeline for Trajectory Prediction
@@
The prediction part of the method heavily relies on a UNet diffusion model with multi-head self-attention layers between blocks. The self-attention layers help to relate the visual memory encoding to different parts of the trajectory, thus facilitating a deeper understanding of how humans interact with environmental features.
The prediction is conditioned on a single 64-dimensional encoded visual memory, and full 100 steps of 6d pose trajectory at 20 Hz. We compress the Visual Memory (VM) from 1x180x360x5 (R-G-B-D-Segmentation) to a 1x64 dimensional embedding using a pre-trained VAE (Encoder). State trajectories are only normalized (100x9) since they are already small. The 9 channels represent x-y-z positions and ortho6d[38] orientation.
We utilize the full 100 steps as we believe in some cases the ego-centric trajectory prediction task can violate pure Markovian assumption. For example, a different past trajectory could likely affect the visibility of the scene in visual memory, causing different exploration preferences in trajectory selection. To verify the effectiveness, we compared it to a pure Markovian assumption in the ablation section.
The UNet takes VM embeddings and normalized states as conditions (Fig. 5). The conditions are integrated by classifier-free guidance, removing the input conditions 10% of the training time. We apply both conditions to all down, up, and self-attention blocks, transforming the shape through a linear layer and adding to the output activation.
The output of the diffusion model is a denoised sequence that contains 100 steps of both normalized ego-centric trajectory and visual memory encoding, resulting in a (100, 9+64) sized prediction for all 100 steps. The normalized trajectory prediction is the first 9, and the embedded VM prediction is the rest. We use the same pre-trained VAE (Decoder) to decode the full 100x180x360x5 visual memory. VM and trajectory are treated as the same latent vector and denoised once, while VAE processes only the visual information.
@@
The visual memory encoding can be decoded to show how visual memory is expected to march forward by the model. Including the visual memory encoding in the diffusion process also provides a foundation for the model to base the predicted trajectory. In practice, we notice that the trajectory prediction samples are more stable with the visual memory prediction. In inference time, the model will take in the same condition and generate a batch of 15 different samples, forming a discrete distribution of the expected future ego trajectory modes.
@@ III-E Hybrid Generation
III-F Diffusion Model Detail
Figure 5: Diffusion Model: Architecture and hybrid generation details. The black and white stripes represent the time-stepping schemes of the inference. In default mode, all steps are diffused sequentially. In our hybrid mode, first perform stripped DDIM steps from 1000 to 0, then DDPM steps from n to 0. Conditions include compressed VM and past states are passed to the embedding layer and concatenated to the block output.
Fig 5 illustrates the detailed model architecture. We use the transformer encoder blocks which contain multi-head self-attention layers. They are stacked between the Down and Up blocks of the UNet. The conditions of the generation (past trajectory and most recent visual memory) are first turned into embeddings and then passed into the Down and Up blocks.
In training, instead of cutting up the dataset into unique sub-trajectories, we take overlapping sub-trajectories to further augment the dataset. This way, there is a total of more than 220,000 diverse sub-trajectories for the model to train on.
At inference time, one can either use DDIM, DDPM, or our hybrid inference method. Specifically, the hybrid inference involves first getting a rough convergence with 20 steps of DDIM, the time steps are sub-sampled uniformly from the total steps. Then the last 5 steps are performed with DDPM steps 5 to 0 to bring the sample to full convergence. The variance we used in DDPM is fixed small, or mathematically:
$x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}z_{\theta}(x_{t},t))+\sigma_{t}z$ (2)
where:
$\sigma_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\beta_{t}$ (3)
IV Egocentric Navigation Dataset